# PoreTrack3D: A Benchmark for Dynamic 3D Gaussian Splatting in Pore-Scale Facial Trajectory Tracking

### v1

@@ Abstract
We introduce PoreTrack3D, the first benchmark for dynamic 3D Gaussian splatting in pore-scale, non-rigid 3D facial trajectory tracking. It contains over 440,000 facial trajectories in total, among which more than 52,000 are longer than 10 frames, including 68 manually reviewed trajectories that span the entire 150 frames. To the best of our knowledge, PoreTrack3D is the first benchmark dataset to capture both traditional facial landmarks and pore-scale keypoints trajectory, advancing the study of fine-grained facial expressions through the analysis of subtle skin-surface motion. We systematically evaluate state-of-the-art dynamic 3D Gaussian splatting methods on PoreTrack3D, establishing the first performance baseline in this domain. Overall, the pipeline developed for this benchmark dataset’s creation establishes a new framework for high-fidelity facial motion capture and dynamic 3D reconstruction. Our dataset are publicly available at: https://github.com/JHXion9/PoreTrack3D
@@ I Introduction
Figure 1: Representative Samples and Trajectory Visualizations of PoreTrack3D. To facilitate visual inspection, the trajectories of different types are color-coded: green denotes the primary trajectories that are successfully tracked across time, blue indicates the supplementary trajectories obtained through temporal completion, and red represents supporting trajectories derived from neighboring keypoints. For clarity in visualization, only the trajectories over the previous 10 frames (out of the total 150) are shown for each displayed frame, and subsequent visualizations follow the same convention.
@@
Recognizing this limitation, recent efforts have extended tracking into 3D. Early attempts such as scene flow [vedula1999three] generalized optical flow to three dimensions but were restricted to instantaneous motion without assessing long-term, occlusion-aware associations. More recent datasets, including TAP-3D [koppula2024tapvid] and Stereo4D [jin2024stereo4d], have begun to define evaluation standards for general 3D motion. However, these benchmarks primarily focus on coarse, keypoint-level trajectories of rigid or semi-rigid objects, and thus fall short of capturing the fine-scale, textureless deformations observed in highly non-rigid surfaces. Their spatial resolution and motion precision—typically at the centimeter level—are insufficient to evaluate methods like Dynamic 3D Gaussian Splatting methods, which aim to model dense, continuous, and sub-millimeter geometric motion. Recent studies [luiten2024dynamic, duisterhof2023deformgs, zhang2024dynamic] have further demonstrated that Dynamic 3D Gaussian Splatting methods achieve superior temporal coherence and tracking accuracy compared with correspondence-based tracking methods—such as PIPS++ [zheng2023pointodyssey], CoTracker [karaev2024cotracker], and SpatialTracker [xiao2024spatialtracker]—by explicitly modeling scene geometry and motion within a unified 3D representation. This geometric consistency allows them to better capture continuous deformations and subtle surface motions, underscoring the need for a dedicated benchmark tailored to such fine-grained, non-rigid dynamics.
@@
Among such non-rigid objects, the human face presents a particularly challenging case. Facial dynamics are driven by complex muscle activations that induce subtle yet critical deformations—such as the protrusion of lips, the strain of skin during a smile, or the emergence of fine wrinkles. Parametric 3D morphable models (3DMMs) [blanz2023morphable], such as FLAME [li2017learning], have provided a compact representation of facial geometry and expressions, and are widely used to synthesize motion data in controlled settings. However, datasets generated from such parametric priors—like FacialFlowNet [lu2024facialflownet]—rely on synthetic renderings with non-realistic textures and do not provide explicit 3D trajectories. As a result, they fail to reflect the true physical behavior of facial surfaces at fine spatial scales, leaving the dense, pore-level motion of real human faces largely unexplored.
@@
To fill this gap, we introduce PoreTrack3D, the first pore-level 3D facial point trajectory benchmark dataset designed for high-precision facial dynamics analysis. Representative samples and trajectory visualizations are shown in Fig. 1. Our main contributions are summarized as follows:
@@
  • PoreTrack3D is introduced as the first benchmark for evaluating dynamic 3D Gaussian splatting methods in pore-scale, non-rigid facial trajectory tracking. In contrast to existing 2D or coarse 3D tracking benchmarks, PoreTrack3D captures both macroscopic landmarks and microscopic surface details, enabling quantitative assessment of subtle, non-rigid facial dynamics.
@@
  • A comprehensive data creation framework is developed, integrating automated trajectory extraction with interactive manual review. This framework supports the generation of high-precision 3D trajectories (up to 150 frames) while maintaining both accuracy and scalability, and is readily extensible to future dynamic facial motion datasets.
@@ II-A Motion Data
In spite of recent progress, existing datasets share a key limitation: they primarily capture large-scale motion. While effective for centimeter-level movements of bodies and objects, this resolution is insufficient for the human face. Facial expressions arise from sub-millimeter skin deformations driven by underlying muscles. The absence of datasets that capture such minute dynamics remains a critical bottleneck for facial analysis research, hindering progress and the development of more precise algorithms. We overcome this limitation by introducing PoreTrack3D, a high-resolution facial dataset with millimeter-level 3D pore-scale trajectories that support precise analysis of fine facial dynamics. Table I reports only datasets that provide ground-truth 3D trajectories, as this requirement is fundamental for a direct and fair comparison; works lacking 3D supervision are therefore omitted.
@@ II-B Dynamic 3D Gaussian Splatting Methods for Tracking and Reconstruction
Dynamic 3D Gaussian Splatting methods have recently emerged as a powerful framework for joint scene reconstruction and motion tracking[luiten2024dynamic, duisterhof2023deformgs, zhang2024dynamic, kerbl20233d]. Compared with traditional NeRF[mildenhall2021nerf] or mesh-based dynamic reconstruction methods[grassal2022neural, li2009robust], Dynamic 3D Gaussian Splatting methods achieve superior rendering fidelity and enables dense, temporally consistent tracking of surface points in real time. By explicitly modeling scene geometry as collections of anisotropic Gaussian primitives, these methods inherently support continuous motion representation and sub-millimeter trajectory estimation—capabilities that are challenging for conventional volumetric or parametric approaches. Recent evaluations [luiten2024dynamic, duisterhof2023deformgs, zhang2024dynamic] further indicate that Dynamic 3D Gaussian Splatting methods surpass correspondence-based tracking methods in temporal coherence and motion accuracy, reinforcing their suitability as the primary focus of our benchmark.
@@ II-C Face-Specific Dynamic 3D Gaussian Splatting Methods for Reconstruction and Tracking
FastAvatar [wu2025fastavatar], GPAvatar [feng2025gpavatar], and Gaussian Déjà-vu [Yan_2025_WACV] further improve the efficiency and fidelity of 3D Gaussian-based facial reconstruction. These methods replace per-frame Gaussian optimization with a neural canonical-space representation, where a network—typically an MLP, Transformer, or GNN—predicts a shared set of Gaussians from multi-view or multi-frame observations. By aggregating information into a unified canonical field, they achieve fast and high-quality avatar generation. However, since the Gaussian identities are not preserved across frames, these canonical-space or graph-based representations lack explicit temporal correspondence and thus cannot provide point-level tracking or fine-grained motion analysis.
Despite these advances, existing face-specific 3D Gaussian Splatting methods mainly focus on improving reconstruction fidelity and visual realism, while neglecting temporal geometric consistency. None of these approaches explicitly establish or evaluate point-level correspondences across time, making it difficult to quantify subtle facial dynamics such as micro-expressions, skin stretching, or pore-level motion. Moreover, the absence of a standardized benchmark prevents fair and reproducible evaluation of temporal accuracy at sub-millimeter precision. To address this gap, we propose PoreTrack3D, a benchmark specifically designed for high-precision facial dynamic analysis, offering millimeter-level 3D trajectories of high-resolution faces for evaluating fine-grained motion and tracking consistency in dynamic 3D Gaussian representations.
@@ III PoreTrack3D
Figure 2: Overview of the proposed four-stage multi-view trajectory reconstruction framework: (A) Mesh reconstruction and 2D-to-3D correspondence mapping, (B) Multi-view patch initialization and tracking, (C) Fine-scale keypoint localization and temporal matching, and (D) Cross-view trajectory integration and temporal completion.
@@
Leveraging its rich capture setting, a four-stage multi-view trajectory reconstruction framework (Fig. 2) is designed to reconstruct and track subtle, non-rigid surface trajectories across frames. The framework consists of four stages: (A) Mesh reconstruction and 2D-to-3D correspondence mapping, (B) Multi-view patch initialization and tracking, (C) Fine-scale keypoint localization and temporal matching, and (D) Cross-view trajectory integration and temporal completion. The technical details of each stage are elaborated below. Note that the 2D-to-3D correspondence mapping is defined in stage (A) and subsequently applied throughout stages (B)–(D).
@@ III-A Mesh Reconstruction and 2D-to-3D correspondence Mapping
To obtain accurate geometric priors for subsequent 2D-to-3D correspondence mapping, we reconstruct a high-fidelity facial mesh by refining the standard COLMAP[schoenberger2016sfm, schoenberger2016mvs] pipeline. Specifically, we employ the PSIFT algorithm[li2015design] to detect dense and stable keypoints, extract discriminative feature descriptors using Pore-HyNet[li2025porescale], and apply GMS matching[bian2017gms] for robust correspondence filtering. The resulting sparse correspondences are then used to estimate camera poses and generate dense point clouds, from which a high-quality facial mesh is obtained through Poisson reconstruction.
@@ III-B Multi-View Patch Initialization and Tracking
These projected centers serve as initialization points for patch tracking. In each view, we crop a $100\times 100$ pixel patch around the projected center and employ the CoTracker [karaev2024cotracker] to track its motion across consecutive frames. This process provides coarse but stable motion tracking for each view, defining a temporally consistent local tracking window where fine-scale pore localization and feature matching are subsequently performed.
III-C Fine-Scale Keypoint Localization and Temporal Matching
Within each tracked patch obtained from the previous step, we perform fine-scale analysis to extract dense and temporally consistent pore-level correspondences. This process involves three components:
@@ III-C Fine-Scale Keypoint Localization and Temporal Matching
Descriptor Extraction. For each keypoint $p_{i_{t}}$, a $64\times 64$ local patch is cropped, resized to $32\times 32$, and processed by the Pore-HyNet network to generate discriminative feature descriptors $d_{i_{t}}$.
@@
Unlike the reference view, the neighboring views do not require keypoint detection. Instead, for each reprojected keypoint, a local patch is extracted to compute descriptors $(d_{\text{neigh1}})_{i_{t}}$} and $(d_{\text{neigh2}})_{i_{t}}$} using Pore-HyNet. The same temporal matching procedure is then applied to these descriptors to enforce geometric and temporal consistency across views.
@@ III-D Cross-View Trajectory Integration and Temporal Completion
Specifically, we start by completing the primary trajectories. For pore-scale keypoints, the primary trajectories correspond to the longest (usually >>70 frames) and most reliable motion sequences. For facial landmarks, the primary trajectories are defined by the landmark points detected on the first frame using PIPNet[jin2021pixel], together with their associated trajectories across the sequence. Suppose a point $p_{i_{t}}$ in a primary trajectory loses its correspondence between frame $t$ and t+1t{+}1. To estimate its motion, we first select its five nearest neighboring keypoints that successfully maintain correspondences between frame $t$ and t+1t{+}1, forming a locally reliable motion reference. The tentative position $p_{i_{t+1}}^{\prime}$ is then predicted by computing a weighted average of their motion vectors, where the weight of each neighbor is inversely proportional to its Euclidean distance from the target point. We subsequently search within a 1-pixel radius of the predicted position $p_{i_{t+1}}^{\prime}$ for any existing keypoint at frame t+1t{+}1. If such a keypoint $p_{i_{t+1}}$ is found, it replaces the predicted position and the two trajectories are merged and treated as a single continuous sequence; otherwise, the predicted position $p_{i_{t+1}}^{\prime}$ is used to complete the missing segment of the original trajectory.
@@ III-E Manual Review
As shown in Fig. 4, the reconstructed keypoint trajectories are projected onto the texture-enhanced reference view to facilitate intuitive inspection of motion trends and spatial coherence. The enhancement is achieved through a GMM-based texture refinement pipeline [pun2014skin], which detects facial landmarks, segments skin regions, and selectively amplifies fine-grained surface details to improve the visibility and contrast of motion traces. Trajectory colors encode their semantic roles: green denotes valid segments of the primary trajectories, blue indicates supplementary trajectories used to fill temporal gaps, and red represents supporting trajectories derived from neighboring keypoints. A color gradient from black to the representative hue (green, blue, or red) indicates temporal direction, with darker colors corresponding to earlier frames. During manual review, only green and blue trajectories are displayed to reduce visual clutter; the red supporting trajectories are included in the paper figures solely for illustration, providing spatial context and aiding in identifying potential inconsistencies.

### v2

@@ Abstract
We introduce PoreTrack3D, the first benchmark for dynamic 3D Gaussian splatting in pore-scale, non-rigid 3D facial trajectory tracking. It contains over 440,000 facial trajectories in total, among which more than 52,000 are longer than 10 frames, including 68 manually reviewed trajectories that span the entire 150 frames. To the best of our knowledge, PoreTrack3D is the first benchmark dataset to capture both traditional facial landmarks and pore-scale keypoints trajectory, advancing the study of subtle facial expressions through the analysis of subtle skin-surface motion. We systematically evaluate state-of-the-art dynamic 3D Gaussian splatting methods on PoreTrack3D, establishing the first performance baseline in this domain. Overall, the pipeline developed for this benchmark dataset’s creation establishes a new framework for high-fidelity facial motion capture and dynamic 3D reconstruction. Our dataset are publicly available at: https://github.com/JHXion9/PoreTrack3D
@@ I Introduction
Figure 1: Representative Samples and Trajectory Visualizations of PoreTrack3D. For each subject, the locally magnified views are arranged in a counterclockwise order starting from the forehead region. To facilitate visual inspection, the trajectories of different types are color-coded: green denotes the primary trajectories that are successfully tracked across time, blue indicates the completion-derived trajectories obtained through temporal completion, and red represents supporting trajectories derived from neighboring keypoints. The direction of each trajectory is indicated by its color gradient, transitioning from black to the corresponding color (red, green, or blue). For clarity in visualization, only the trajectories over the previous 10 frames (out of the total 150) are shown for each displayed frame, and subsequent visualizations follow the same convention.
@@
Recognizing this limitation, recent efforts have extended tracking into 3D. Early attempts such as scene flow [vedula1999three] generalized optical flow to three dimensions but were restricted to instantaneous motion without assessing long-term, occlusion-aware associations. More recent datasets, including TAP-3D [koppula2024tapvid] and Stereo4D [jin2024stereo4d], have begun to define evaluation standards for general 3D motion. However, these benchmarks primarily focus on coarse, keypoint-level trajectories of rigid or semi-rigid objects, and thus fall short of capturing the subtle, textureless deformations observed in highly non-rigid surfaces. Their spatial resolution and motion precision—typically at the centimeter level—are insufficient to evaluate methods like Dynamic 3D Gaussian Splatting methods, which aim to model dense, continuous, and sub-millimeter geometric motion. Recent studies [luiten2024dynamic, duisterhof2023deformgs, zhang2024dynamic] have further demonstrated that Dynamic 3D Gaussian Splatting methods achieve superior temporal coherence and tracking accuracy compared with correspondence-based tracking methods—such as PIPS++ [zheng2023pointodyssey], CoTracker [karaev2024cotracker], and SpatialTracker [xiao2024spatialtracker]—by explicitly modeling scene geometry and motion within a unified 3D representation. This geometric consistency allows them to better capture continuous deformations and subtle surface motions, underscoring the need for a dedicated benchmark tailored to such subtle, non-rigid dynamics.
@@
Among such non-rigid objects, the human face presents a particularly challenging case. Facial dynamics are driven by complex muscle activations that induce subtle yet critical deformations—such as the protrusion of lips, the strain of skin during a smile, or the emergence of fine wrinkles. Parametric 3D morphable models (3DMMs) [blanz2023morphable], such as FLAME [li2017learning], have provided a compact representation of facial geometry and expressions, and are widely used to synthesize motion data in controlled settings. However, datasets generated from such parametric priors—like FacialFlowNet [lu2024facialflownet]—rely on synthetic renderings with non-realistic textures and do not provide explicit 3D trajectories. As a result, they fail to reflect the true physical behavior of facial surfaces at fine spatial scales, leaving the dense, pore-scale motion of real human faces largely unexplored.
@@
To fill this gap, we introduce PoreTrack3D, the first pore-scale 3D facial point trajectory benchmark dataset designed for high-precision facial dynamics analysis. Representative samples and trajectory visualizations are shown in Fig. 1. Our main contributions are summarized as follows:
@@
  • PoreTrack3D is introduced as the first benchmark for evaluating dynamic 3D Gaussian splatting methods in pore-scale, non-rigid facial trajectory tracking. In contrast to existing 2D or coarse 3D tracking benchmarks, PoreTrack3D captures both facial landmark trajectories and pore-scale trajectories, enabling quantitative assessment of subtle, non-rigid facial dynamics.
@@
  • A comprehensive data creation framework is developed, integrating automated trajectory extraction with interactive manual review. This framework supports the generation of high-precision 3D trajectories while maintaining both accuracy and scalability.
@@ II-A Motion Data
In spite of recent progress, existing datasets share a key limitation: they primarily capture large-scale motion. While effective for centimeter-level movements of bodies and objects, this resolution is insufficient for the human face. Facial expressions arise from millimeter skin deformations driven by underlying muscles. The absence of datasets that capture such subtle dynamics remains a critical bottleneck for facial analysis research, hindering progress and the development of more precise algorithms. We overcome this limitation by introducing PoreTrack3D, a high-resolution facial dataset with millimeter-level 3D pore-scale trajectories that support precise analysis of fine facial dynamics. Table I reports only datasets that provide 3D trajectories, as this requirement is fundamental for a direct and fair comparison; works lacking 3D supervision are therefore omitted.
@@ II-B Dynamic 3D Gaussian Splatting Methods for Tracking and Reconstruction
Dynamic 3D Gaussian Splatting methods have recently emerged as a powerful framework for joint scene reconstruction and motion tracking[luiten2024dynamic, duisterhof2023deformgs, zhang2024dynamic, kerbl20233d]. Compared with traditional NeRF[mildenhall2021nerf] or mesh-based dynamic reconstruction methods[grassal2022neural, li2009robust], Dynamic 3D Gaussian Splatting methods achieve superior rendering fidelity and enables dense, temporally consistent tracking of surface points in real time. By explicitly modeling scene geometry as collections of anisotropic Gaussian primitives, these methods inherently support continuous motion representation and millimeter trajectory estimation—capabilities that are challenging for conventional volumetric or parametric approaches. Recent evaluations [luiten2024dynamic, duisterhof2023deformgs, zhang2024dynamic] further indicate that Dynamic 3D Gaussian Splatting methods surpass correspondence-based tracking methods in temporal coherence and motion accuracy, reinforcing their suitability as the primary focus of our benchmark.
@@ II-C Face-Specific Dynamic 3D Gaussian Splatting Methods for Reconstruction and Tracking
FastAvatar [wu2025fastavatar], GPAvatar [feng2025gpavatar], and Gaussian Déjà-vu [Yan_2025_WACV] further improve the efficiency and fidelity of 3D Gaussian-based facial reconstruction. These methods replace per-frame Gaussian optimization with a neural canonical-space representation, where a network—typically an MLP, Transformer, or GNN—predicts a shared set of Gaussians from multi-view or multi-frame observations. By aggregating information into a unified canonical field, they achieve fast and high-quality avatar generation. However, since the Gaussian identities are not preserved across frames, these canonical-space or graph-based representations lack explicit temporal correspondence and thus cannot provide point-level tracking or subtle motion analysis.
Despite these advances, existing face-specific 3D Gaussian Splatting methods mainly focus on improving reconstruction fidelity and visual realism, while neglecting temporal geometric consistency. None of these approaches explicitly establish or evaluate point-level correspondences across time, making it difficult to quantify subtle facial dynamics such as micro-expressions, skin stretching, or pore-scale motion. Moreover, the absence of a standardized benchmark prevents fair and reproducible evaluation of temporal accuracy at millimeter precision. To address this gap, we propose PoreTrack3D, a benchmark specifically designed for high-precision facial dynamic analysis, offering millimeter-level 3D trajectories of high-resolution faces for evaluating subtle motion and tracking consistency in dynamic 3D Gaussian representations.
@@ III PoreTrack3D
Figure 2: Overview of the proposed four-stage multi-view trajectory reconstruction framework: (A) Mesh reconstruction and 2D-to-3D correspondence mapping, (B) Multi-view patch initialization and tracking, (C) Pore-scale keypoint localization and temporal matching, and (D) Cross-view trajectory integration and temporal completion.
@@
Leveraging its rich capture setting, a four-stage multi-view trajectory reconstruction framework (Fig. 2) is designed to reconstruct and track subtle, non-rigid surface trajectories across frames. The framework consists of four stages: (A) Mesh reconstruction and 2D-to-3D correspondence mapping, (B) Multi-view patch initialization and tracking, (C) Pore-scale keypoint localization and temporal matching, and (D) Cross-view trajectory integration and temporal completion. The technical details of each stage are elaborated below. Note that the 2D-to-3D correspondence mapping is defined in stage (A) and subsequently applied throughout stages (B)–(D).
@@ III-A Mesh Reconstruction and 2D-to-3D correspondence Mapping
To obtain accurate geometric priors for subsequent 2D-to-3D correspondence mapping, we reconstruct a high-fidelity facial mesh by refining the standard COLMAP[schoenberger2016sfm, schoenberger2016mvs] pipeline. Specifically, we employ the PSIFT algorithm[li2015design] to detect dense and stable keypoints, extract discriminative feature descriptors using Pore-HyNet-V4[li2025porescale], and apply GMS matching[bian2017gms] for robust correspondence filtering. The resulting sparse correspondences are then used to estimate camera poses and generate dense point clouds, from which a high-quality facial mesh is obtained through Poisson reconstruction.
@@ III-B Multi-View Patch Initialization and Tracking
These projected centers serve as initialization points for patch tracking. In each view, we crop a $100\times 100$ pixel patch around the projected center and employ the CoTracker [karaev2024cotracker] to track its motion across consecutive frames. This process provides coarse but stable motion tracking for each view, defining a temporally consistent local tracking window where pore-scale keypoints localization and feature matching are subsequently performed.
III-C Pore-Scale Keypoint Localization and Temporal Matching
Within each tracked patch obtained from the previous step, we perform pore-scale analysis to extract dense and temporally consistent correspondences. This process involves three components:
@@ III-C Pore-Scale Keypoint Localization and Temporal Matching
Descriptor Extraction. For each keypoint $p_{i_{t}}$, a $64\times 64$ local patch is cropped, resized to $32\times 32$, and processed by the Pore-HyNet-V4 network to generate discriminative feature descriptors $d_{i_{t}}$.
@@
Unlike the reference view, the neighboring views do not require keypoint detection. Instead, for each reprojected keypoint, a local patch is extracted to compute descriptors $(d_{\text{neigh1}})_{i_{t}}$} and $(d_{\text{neigh2}})_{i_{t}}$} using Pore-HyNet-V4. The same temporal matching procedure is then applied to these descriptors to enforce geometric and temporal consistency across views.
@@ III-D Cross-View Trajectory Integration and Temporal Completion
Specifically, we start by completing the primary trajectories. For pore-scale keypoints, the primary trajectories correspond to the longest (usually >>70 frames) and most reliable motion sequences. For facial landmarks, the primary trajectories are defined by the landmark points detected on the first frame using PIPNet[jin2021pixel], together with their associated trajectories across the sequence. Suppose a point $p_{i_{t}}$ in a primary trajectory loses its correspondence between frame $t$ and t+1t{+}1. To estimate its motion, we first select its five nearest neighboring keypoints that successfully maintain correspondences between frame $t$ and t+1t{+}1, forming a locally reliable motion reference. The tentative position $p_{i_{t+1}}^{\prime}$ is then predicted by computing a weighted average of their motion vectors, where the weight of each neighbor is inversely proportional to its Euclidean distance from the target point. We subsequently search within a 1-pixel radius of the predicted position $p_{i_{t+1}}^{\prime}$ for any existing keypoint at frame t+1t{+}1. If such a keypoint $p_{i_{t+1}}$ is found, it replaces the predicted position and the two trajectories are merged and treated as a single continuous sequence; otherwise, the predicted position $p_{i_{t+1}}^{\prime}$ is used to complete the missing segment of the original trajectory, which we refer to as completion-derived trajectories.
@@ III-E Manual Review
As shown in Fig. 4, the reconstructed keypoint trajectories are projected onto the texture-enhanced reference view to facilitate intuitive inspection of motion trends and spatial coherence. The enhancement is achieved through a GMM-based texture refinement pipeline [pun2014skin], which detects facial landmarks, segments skin regions, and selectively amplifies subtle surface details to improve the visibility and contrast of motion traces. Trajectory colors encode their semantic roles: green denotes valid segments of the primary trajectories, blue indicates completion-derived trajectories used to fill temporal gaps, and red represents supporting trajectories derived from neighboring keypoints. A color gradient from black to the representative hue (green, blue, or red) indicates temporal direction, with darker colors corresponding to earlier frames. During manual review, only the green hollow circles are displayed to reduce visual clutter; the red supporting trajectories are included in the paper figures solely for illustration, providing spatial context and aiding in identifying potential inconsistencies.